OpenStudy (anonymous):

solve the proportion: 3/x=(x+7)/(5x+3) a. x= -1and 9 b. x= 1 and 9 c. x= -1 and -9 d. x= 1 and -9

OpenStudy (anonymous):

\[\frac{3}{x}=\frac{x+7}{5x+3}\] solve \[3(5x+3)=x(x+7)\]

OpenStudy (anonymous):

you will get a quadratic equation which is why there are two solutions

OpenStudy (anonymous):

@satellite73 should i have gotten 15x +9 = x^2 +7x

OpenStudy (anonymous):

yes as a first step

OpenStudy (anonymous):

the set it equal to zero and solve by factoring someone went to a lot of trouble to cook a problem up so that the quadratic would factor the least you can do is not solve using a real technique like the quadratic equation, just factor it